An asset has an average return of 10.15 percent and a standard deviation of 19.05 percent. What range of returns should you expect to see with a 68 percent probability? A.−47.00% to 67.30%

B −18.43% to 38.73%

C −8.90% to 29.20%

D −27.95% to 48.25%

E −8.90% to 11.40%
Answer #1

Correct Answer - Option C

Based on 68-95-99 rule,

68% 95% 99% -30 +10 -20 -10 +20 +30

Now, this implies the confidence interval for 68% is +/- 1 * Standard deviation

Upper interval = Mean + 1 * Standard deviation

Upper interval = 10.15% + 1 * 19.05% = 29.20%

Lower interval = Mean - 1 * Standard deviation

Upper interval = 10.15% - 1 * 19.05% = -8.90%

68% 95% 99% -16 +16 +26 +30